Find a number r such that (1 + r 1075 ) (1075) = 7

Using four rectangles, show that area(x 2 , 1, 2) < 2.72.
Answer #1

To find r И. (1791015) (1075) = 7 It a 1075= 7 1675 r 1075 = 7 - 1075 -1068 1075 r = -1068 X 1075 r 1075 1075 r= 1068 (1075)